There are 30% registered voters who support candidate 1. A pollster will obtain a random sample of 2200 registered voters. Find the probability that at most 1500 of the individuals in the sample support 1?
Answer #1

n = 2200, p = 0.3

mean = 2200*0.3 = 660

sd = sqrt(n*p*q) = 21.49 (approx)

P(at most 1500) = [Due to continuoity correction]

So, the probability is 1 in this case.

I hope this will help. Thank you.